# Image Intrinsic Scale Assessment: Bridging the Gap Between Quality and Resolution

### v2

@@ 6.2 Results
We integrate the proposed approach with several state-of-the-art NR-IQA methods retrained for IISA. For a comprehensive analysis, we consider baselines based on diverse strategies. Specifically, we consider approaches that train a network from scratch with supervised learning (DBCNN [34] and TOPIQ [5]), methods that train a regressor head on top of an encoder pre-trained with self-supervised learning (CONTRIQUE [25] and ARNIQA [2]), and VLM-based models that rely on prompt learning [32] (CLIP-IQA+ [32] and QualiCLIP [1]). For each baseline, we train two variants on the IISA-DB dataset: a base one using only the ground-truth labels, and one also leveraging the weak labels generated with WIISA.
@@ 6.3 Discussion
We conduct controlled experiments to assess how key components affect our approach’s performance, namely: 1) the number of weak labels $n_{wl}$; 2) the downscaling threshold $\delta$; 3) the interpolation algorithm used to generate the weak labels. For each experiment, we keep other hyper-parameters fixed unless otherwise stated. We consider only the TOPIQ model for its superior performance among the studied approaches.
Tab. 3 shows the results. Regarding the number of weak labels, the experiments show that increasing the value of $n_{wl}$ leads to better results up to $n_{wl}=2$, but beyond that performance declines. We hypothesize that too many weak labels introduce overly redundant information, as the image content remains constant across them, with only the scale varying. Likewise, a high downscaling threshold leads to the weak labels being overly similar to the ground-truth ones, thus reducing the performance improvement. Conversely, lower $\delta$ values allow the weakly labeled IIS to be higher (see Eq. 2), possibly resulting in them being out of the ground-truth distribution (the average IIS in IISA-DB is 0.347). Finally, we observe that using the bilinear and bicubic interpolation algorithms to generate the weak labels leads to marginally lower performance than using the Lanczos one. This finding suggests that the interpolation algorithm has limited impact in training models with WIISA.
@@ Acknowledgements
Given the formulation of IIS reported in Eq. 1, we can employ the quality scores predicted by a pre-trained NR-IQA method to estimate the IIS automatically. Specifically, given $n_{s}$ uniformly sampled scales $s$ in the range $[s_{lb},1]$, we can use a pre-trained NR-IQA model to assess the quality of each downscaled version $I^{s}$ of an input image $I$ and then find the scale for which the predicted quality is the highest. This would be an estimate for the IIS of image $I$.
Following the evaluation protocol described in Sec. 6.1, we assess the performance of this zero-shot multi-scale approach on the proposed IISA-DB dataset. We employ two versions of the TOPIQ [5] model pre-trained on the SPAQ [9] and UHD-IQA [15] datasets, which feature high-resolution images similar to those in IISA-DB. We use $n_{s}=100$ scales and report the results in Tab. S1. We observe that the zero-shot multi-scale approach achieves unsatisfactory performance, regardless of the pre-training dataset. We attribute this to NR-IQA models struggling to handle the change in perceptual quality caused by downscaling, as noted in [33, huang2024high]. In addition, the multi-scale approach requires multiple model forward passes to obtain a single IIS prediction, which can be inefficient.

### v3

@@ 6.2 Results
We integrate the proposed approach with several state-of-the-art NR-IQA methods retrained for IISA. For a comprehensive analysis, we consider baselines based on diverse strategies. Specifically, we consider approaches that train a network from scratch with supervised learning (DBCNN [35] and TOPIQ [5]), methods that train a regressor head on top of an encoder pre-trained with self-supervised learning (CONTRIQUE [26] and ARNIQA [2]), and VLM-based models that rely on prompt learning [33] (CLIP-IQA+ [33] and QualiCLIP+ [1]). For each baseline, we train two variants on the IISA-DB dataset: a base with only ground-truth labels, and another that also uses weak labels from WIISA.
@@ 6.3 Discussion
We conduct controlled experiments to assess how key components affect our approach’s performance, namely: 1) the number of weak labels $n_{wl}$; 2) the downscaling threshold $\delta$; 3) the interpolation algorithm used to generate the weak labels. For each experiment, we keep other hyperparameters fixed unless otherwise stated. We consider only the TOPIQ model for its superior performance among the studied approaches.
Tab. 3 shows the results. Regarding the number of weak labels, the experiments show that increasing the value of $n_{wl}$ leads to better results up to $n_{wl}=2$, but beyond that performance declines. We hypothesize that too many weak labels introduce overly redundant information, as the image content remains constant across them, with only the scale varying. Likewise, a high downscaling threshold leads to the weak labels being overly similar to the ground-truth ones, thus reducing the performance improvement. Conversely, lower $\delta$ values allow the weakly labeled IIS to be higher (see Eq. 2), possibly resulting in them being out of the ground-truth distribution (the average IIS in IISA-DB is 0.347). Finally, we observe that using the bilinear and bicubic interpolation algorithms to generate the weak labels leads to marginally lower performance than using the Lanczos one. This finding suggests that the interpolation algorithm has a limited impact in training models with WIISA.
@@ Acknowledgements
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure S6: Visualization of TOPIQ’s model attention.
@@
S2.5 Examples of Attention Maps
We visualize the attention maps of the TOPIQ model trained with our WIISA approach in Fig. S6, using the method described in the original paper [5]. Fig. S6 (a) shows an image featuring a small foreground object in focus against a blurry background, while Fig. S6 (b) depicts a high-semantic object surrounded by high-frequency texture. The model attends to in-focus and high-semantic regions, suggesting that IIS predictions are primarily driven by the degradation of high-level semantic content.
@@
Given the formulation of IIS reported in Eq. 1, we can employ the quality scores predicted by a pre-trained NR-IQA method to estimate the IIS automatically. Specifically, given $n_{s}$ uniformly sampled scales $s$ in the range $s_{lb},1]$, we can use a pre-trained NR-IQA model to assess the quality of each downscaled version $I^{s}$ of an input image $I$ and then find the scale for which the predicted quality is the highest. This would be an estimate for the IIS of image $I$.
Following the evaluation protocol described in Sec. 6.1, we assess the performance of this zero-shot multi-scale approach on the proposed IISA-DB dataset. We employ two versions of the TOPIQ [5] model pre-trained on the SPAQ [9] and UHD-IQA [15] datasets, which feature high-resolution images similar to those in IISA-DB. We use $n_{s}=100$ scales and report the results in Tab. S1. We observe that the zero-shot multi-scale approach achieves unsatisfactory performance, regardless of the pre-training dataset. We attribute this to NR-IQA models struggling to handle the change in perceptual quality caused by downscaling, as noted in [34, 16]. In addition, the multi-scale approach requires multiple model forward passes to obtain a single IIS prediction, which can be inefficient.
@@
The scope of this study is focused on the impact of low-level distortions in User-Generated Content (UGC) images. Consequently, the applicability of our findings to specialized domains governed by different perceptual criteria, such as text-heavy images where readability is the primary quality indicator, remains largely unexplored. Extending the proposed framework to these areas is a promising direction for future work, which would involve adapting protocols to capture domain-specific quality factors.